OpenStudy (anonymous):

I'm not sure why I'm getting this wrong. For the sets A= {1, 2, . . . , 10} and B = {2, 4, 6, 9, 12, 25}, consider the statements P: A is a subset of B. Q: |A-B| = 6. Is P V Q true or false? I said it's false because neither is true. A is not a subset of B, and the difference of the cardinalities of the sets is not 6. However, my book said it's true. I'm confused. :(

OpenStudy (jamesj):

What is the set A - B? It is \[A - B = \{ a \in A \ | \ a \notin B \} = \{ 1, 3, 5, 7, 8, 10 \}\]

OpenStudy (jamesj):

I'm sure you now see your problem.

jimthompson5910 (jim_thompson5910):

P: A is a subset of B This is false since 1 (the first element in set A) is NOT in set B --------------------------- Q: |A-B| = 6 To find A - B, start with A and erase all elements found in set B A = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10} A-B = {1, 3, 5, 7, 8, 10} Now count the number of elements. There are 6 elements here, so |A - B| = 6 So Q: |A-B| = 6 is true Therefore, P v Q is true

OpenStudy (anonymous):

Thanks Jim. I didn't understand how that cardinality was working there. It makes sense now!

OpenStudy (anonymous):

Thanks James, yes. Awesome.

OpenStudy (jamesj):

In the notation of your last question ||A - B|| = ||A|| - ||A n B|| = 10 - 4 and the critical thing is there are other elements of B which are not in A; but they play no role here. That's where you were getting caught up.

OpenStudy (anonymous):

I thought it was the difference of the cardinalities. My professor and book did not explain this as far as I can tell.

OpenStudy (anonymous):

As this example, show us, no. Here's another example A = { 2 } B = { prime numbers less than 1,000,000,000 } Clearly the cardinality of A - B is zero, but it is not the cardinality of A (which is 1) less the cardinality of B (which is something quite a bit bigger than 1)

jimthompson5910 (jim_thompson5910):

|A - B| = the number of elements in set A - B This is the same as saying that |A| = the number of elements in set A

OpenStudy (anonymous):

Thanks everyone, it is much clearer now. Math logic is new to me.